the perimeter of a television screen is 58 inches . it is 18 inches wide. how tall is it?
Answer by Theo(13342)

perimeter equals 2W + 2L
W equal the width
L equal the length.
since W is equal to 18, then 2W is equal to 36.
2W + 2L = 58
36 + 2L = 58
2L = 58 - 36
2L = 22
L = 11
you get:
W = 18
L = 11
2W + 2L becomes 2*18 + 2*11 becomes 36 + 22 becomes 58.
measurements look good.
the television is 11 inches tall.